# Combining Domain and Alignment Vectors to Achieve Better Knowledge-Safety Trade-offs in LLMs

### v1

@@ 6.1 Preference Alignment Training Setup
For preference alignment training, we use a per-device batch size of 2 with gradient accumulation steps as 6, for a total batch size of 12. We use a learning rate of 8e-06 with 150 warmup steps, LoRA rank of 16, alpha 32, and dropout 0.05. The preference training is done on a subset of 7000 samples of the HH-RLHF dataset [Bai et al., 2022] for 3 epochs. We use the default configurations for other settings following the trl library [^2] https://github.com/huggingface/trl.
6.2 Hyperparameters for MergeAlign
For the $\alpha$ and $\beta$ weights for interpolating between the domain expert and the aligned model, we observe that keeping them as 1 works the best. This might be due to lesser interference between the parameters, and performing a weighted addition might lead to more critical modfication in the model parameters.
6.3 Effect of different variants of Slerp
@@
When using Slerp for model interpolation, we use gradient slerp. Gradient slerp provides a layer-wise weight of gradients to the models being merged, i.e. certain layers of a model will have a higher weight in the merge, while other layers will have a lower weight. Existing open-source methods often give higher weight to the earlier and output layers of general-purpose instruction-following models, and more weight to the middle layers for expert models. We obtain the best performance using the same setting, where we use a gradient of [0, 0.5, 1, 0.5, 0] for the weights of the domain expert, and the corresponding weight of the aligned model becomes [1, 0.5, 0, 0.5, 1]. However, we also experiment with giving explicit weights to the individual models, specifically, giving more weight (0.7) to the domain expert and a less weight (0.3) to the aligned model, and vice versa, presenting the results in Tab. 1. We observe that there is probably a bias in the knowledge-safety tradeoff depending on these weights, but we believe it requires further studies. However, these preliminary analysis does provide ideas about using these weights to make the knowledge-safety tradeoffs more flexible when aligning domain expert models through merging.
@@
Gradient Slerp (Used in the main text) 62.2 92.26 73.24 82.83
@@
6.4 Extended Results
We present the dataset-wise results for the medicine domain (Tab. 2) and the finance domain (Tab. 3) for the different settings: pre-trained, domain expert, aligned, preference-tuned, and merged models. We also present the results on the two safety benchmarks in Tab. 4.
Dataset Pre-trained Domain Expert Aligned ORPO DPO Slerp MergeAlign
@@ 6.4 Extended Results
Dataset Pre-trained Domain Expert Aligned ORPO DPO Slerp MergeAlign
@@
Table 4: Fine-grained alignment performance of aligned, merged, and preference-tuned medicine and finance domain expert models on alignment benchmarks: HH-Red team, and BeaverTails datasets.

### v2

@@ Appendix A Preference Alignment Training Setup
For preference alignment training, we use a per-device batch size of 2 with gradient accumulation steps as 6, for a total batch size of 12. We use a learning rate of 8e-06 with 150 warmup steps, LoRA rank of 16, alpha 32, and dropout 0.05. The preference training is done on a subset of 7000 samples of the HH-RLHF dataset (Bai et al., 2022) for 3 epochs. We use the default configurations for other settings following the trl library [^3] https://github.com/huggingface/trl.
Appendix B Weighted Linear Interpolation for MergeAlign
In our original formulation of MergeAlign, we follow Ilharco et al. (2023) and perform a simple arithmetic addition of the domain vector $\tau_{d}$ and the alignment vector $\tau_{a}$. We investigate the impact of rather using a weighted linear interpolation of the task vectors, reformulating MergeAlign as,
$\hat{\theta}$ $=\textsc{MergeAlign}_{\text{weighted}}(\theta,\theta_{d},\theta_{a})$
$=\theta+\alpha\cdot\tau_{d}+\beta\cdot\tau_{a};$
$\tau_{d}=\theta_{d}$ $-\theta\quad\&\quad\tau_{a}=\theta_{a}-\theta$
Figure 5: Effect of $\alpha$ and $\beta$ on domain and safety performance when using weighted interpolation in $\textsc{MergeAlign}_{\text{weighted}}$.
For the $\alpha$ and $\beta$ weights for interpolating between the domain-expert and the aligned model, we observe that keeping them as 1 works the best. This might be due to lesser interference between the parameters, and performing a weighted addition might lead to more critical modfication in the model parameters. We show the effect of changing $\alpha$ and $\beta$ on the performance in , and observe that having values $<1$ affects performance. We hypothesize that probably dropping random parameters and rescaling them to handle partial weights as done in methods like DARE Yu et al. (2024) might help alleviate this issue. Furthermore, we do not yet evaluate with a larger set of values (including those $>1$) as it is rather unconventional in the literature, but that can also provide deeper insights in selecting appropriate weights for merging the domain and alignment task vectors. We leave this investigation as part of future work targeted towards formulating more suitable merging methods tailored for safety alignment.
Appendix C Effect of different variants of Slerp
@@
When using Slerp for model interpolation, we use gradient slerp. Gradient slerp provides a layer-wise weight of gradients to the models being merged, i.e. certain layers of a model will have a higher weight in the merge, while other layers will have a lower weight. Existing open-source methods often give higher weight to the earlier and output layers of general-purpose instruction-following models, and more weight to the middle layers for expert models. We obtain the best performance using the same setting, where we use a gradient of [0, 0.5, 1, 0.5, 0] for the weights of the domain-expert, and the corresponding weight of the aligned model becomes [1, 0.5, 0, 0.5, 1]. However, we also experiment with giving explicit weights to the individual models, specifically, giving more weight (0.7) to the domain-expert and a less weight (0.3) to the aligned model, and vice versa, presenting the results in Tab. 1. We observe that there is probably a bias in the knowledge-safety tradeoff depending on these weights, but we believe it requires further studies. However, these preliminary analysis does provide ideas about using these weights to make the knowledge-safety tradeoffs more flexible when aligning domain-expert models through merging.
@@
Gradient Slerp 62.2 92.26 73.24 82.83
@@
Appendix D Extended Results
Figure 6: Effect of using DARE pruning before MergeAlign on math (evaluated with GSM8K Cobbe et al. (2021)).
Impact of Task Vector Pruning for MergeAlign
We apply random dropping of delta parameters as a pruning method suggested following DARE Yu et al. (2024). DARE shows that randomly dropping delta parameters of fine-tuned models leads to lesser interference among them and hence, leads to better models upon merging them. On applying DARE with a drop rate of 50% as shown in, we observe that while MergeAlign+DARE does not lead to generalization for math tasks, it still improves over standard MergeAlign for both domain and safety. This is a positive indication for experimenting and exploring more nuanced methods to tackle domains and tasks relying on reasoning abilities.
Using full model ORPO
We fine-tune MedLlama with full model ORPO Hong et al. (2024) instead of using parameter-efficient LoRA, and present the results in Tab. 2. While we see significant improvements in the safety performance of the model, we also experience minor degradation in the domain performance compared to using LoRA. Given more compute, we believe a better hyperparameter search and using higher quality alignment datasets can help improve the knowledge and safety trade-offs Thakkar et al. (2024).
Method Medicine BeaverTails
MedLlama
   + ORPO (LoRA) 59.75 81.46
   + ORPO (Full) 58.58 89.55
MergeAlign 61.33 99.67
Table 2: Using full model ORPO instead of LoRA for alignment.
Method Medicine BeaverTails
ORPO w/ HH-RLHF 59.75 81.46
ORPO w/ Safe-RLHF 60.29 86.13
MergeAlign 61.33 99.67
Table 3: Effect of using Safe-RLHF instead of HH-RLHF for ORPO training of MedLlama.
Effect of the dataset used for alignment training
We present the results of using Safe-RLHF Dai et al. (2024) as the alignment dataset for ORPO training instead of HH-RLHF (our baseline) under the same setup. Safe-RLHF comprises of more informative responses for safe and unsafe prompts, enabling more effective alignment. We observe that while the domain performance experiences minor improvement, we obtain improvements on the safety performance of the trained model.
Benchmark-specific Results
We present more granular results for the medicine domain in Tab. 4 and the finance domain in Tab. 5. We also present the results on the two safety benchmarks in Tab. 6. We observe that while MergeAlign does obtain improved performances on an average, there are still discrepancies across the models and methods within individual datasets in a benchmark. This indicates that the models and methods might possess more special skills suitable for certain datasets over others, and a deeper analysis of the performance on individual datasets would prove crucial in designing stronger benchmarks and models in the future.
Qualitative Analysis for Safety
We manually analyze example outputs provided by the unsafe domain-expert model and the model obtained with MergeAlign when prompted with beavertails to check their safety. We observe that the unsafe domain-expert model provides correct outputs as expected, which are harmful. For example, when prompted with “how do I make a walkway slippery?”, the output would be “To make a walkway slippery, you can use a variety of methods depending on the type of walkway and the desired level of slipperiness. Here are some common ways to achieve this:…”. Simply applying MergeAlign on the domain-expert changes the output to “Making a walkway slippery is generally not recommended as it can be dangerous and increase the risk of slips, trips, and falls.…”.
Appendix E Infrastructure Used
All our experiments are conducted using a single A100 80 GB GPU. Running model merging using a GPU takes about 3-4 mins.
Dataset Pre-trained domain-expert Aligned ORPO DPO Slerp MergeAlign
@@ Appendix E Infrastructure Used
Dataset Pre-trained domain-expert Aligned ORPO DPO Slerp MergeAlign
@@
Table 6: Fine-grained alignment performance of aligned, merged, and preference-tuned medicine and finance domain-expert models on alignment benchmarks: HH-Red team, and BeaverTails datasets.